OpenStudy (anonymous):

find the values of x between 0 and 2pi where the tangent line to the graph of y=sinxcosx is horizontal.

OpenStudy (rulnick):

Answer is pi/4, 3pi/4, 5pi/4, 7pi/4. Need the details?

OpenStudy (rulnick):

Details: Tangent lines are horizontal where derivative is 0: 0 = dy/dx = -sin^2(x) + cos^2(x) = cos(2x) But cos(x) is 0 at x=pi/2, 3pi/2, 5pi/2, 7pi/2 so cos(2x) is 0 at x=pi/4, 3pi/4, 5pi/4, 7pi/4. That's it.